OpenStudy (anonymous):

simultaneously solve y=-0.05t(t-20) and y=-0.1t(t-28)

OpenStudy (anonymous):

even if you remotely know how to do this. Any help is always welcome.

OpenStudy (anonymous):

@jb1515g could you please help me!!

OpenStudy (anonymous):

Solve for t in one of the equations, then plug that expression into the other equation to solve for y (which will get you a number). Then plug that number in to solve for t's value.

OpenStudy (jack1):

y=-0.05t(t-20) and y=-0.1t(t-28) so the equations can be solved the same way u do normal substitution y = -0.05t(t-20) = t - 0.05t^2 y = -0.1t(t-28) = 2.8t - 0.1t^2 t - 0.05t^2 = 2.8t - 0.1t^2 divide both sides by t 1 - 0.05t = 2.8 - 0.1t therefore t = ...?

Miracrown (miracrown):

@hmaug4 follow Jack's explanation its pretty sound!

OpenStudy (anonymous):

I have a question about the divide everything by t, what does this do? @Miracrown and @Jack1

OpenStudy (jack1):

as everything is a multiple of t, the solution t = 0 is valid (therefore y = 0)

OpenStudy (jack1):

@hmaug4 it proves the y can equal 0 at t = 0, other than that it merely simplifies the equation, in the same way that 400x + 200x^2 = 800x you can divide both sides of the equation by 100 or by x or by both and you'll still arrive at the same answer (x = 2) 4x + 2x^2 = 8x (divide by 100) 400 + 200x = 800 (divide by x) 4 + 2x = 8 (divide by both)

OpenStudy (jack1):

now if the above question was phrased as y = 400x +200x^2 and y = 800x then the solution x = 0 would still be accurate (as y = 0 if x = 0 is TRUE for both statements)

OpenStudy (jack1):

so 2 possible answers, x = 2 and x = 0

OpenStudy (anonymous):

okay thankyou @Jack1 I really appreciate this. Thanks to u I will now be able to pass my maths assignment. Thanks for the explanation!!!!!!!!!!!!!!!

OpenStudy (jack1):

all good dude, yw hey